OpenStudy (anonymous):

Help me find the solutions in the given domain 0 less than and equal to x<2pi for the equation: cos(2x)=1/sqrt2?

OpenStudy (anonymous):

Let me make sure I'm getting this right. The domain is:\[0 \le x \le 2\pi\] Right? Keep in mind that \[\frac{ 1 }{ \sqrt{2} } = \frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

Should be able to solve for x, keeping in mind (man I say that a lot) it's not going to be one of the standard angles on the unit circle since its 2*x

OpenStudy (anonymous):

\[0 \le x < 2 \pi\]

OpenStudy (anonymous):

There should be two solutions to the problem, both when x is positive on the unit circle.

OpenStudy (anonymous):

So, it would be 7pi/4 and pi/4 then, right?

OpenStudy (psymon):

Thats where cos(x) is equal to those values, but you need 2x.

OpenStudy (anonymous):

How would you solve for 2x then?

OpenStudy (anonymous):

Close. But its 2*x. so, at pi/4 in this equation, you're really doing 2*(pi/4) = pi/2

OpenStudy (psymon):

2x = pi/4, 7pi/4 Divide by 2.

OpenStudy (anonymous):

First note:\[\bf \frac{ 1 }{ \sqrt{2} }=\frac{ \sqrt{2} }{ 2 }\]Now in the interval \(\bf 0 \le x < 2 \pi\), cos(2x) hits \(\bf 1/\sqrt{2}\) twice as many times as cos(x). We know that for cosine to be \(\bf 1/\sqrt{2}\), the angle inside the brackets must be in the form \(\bf \pi/2 + 2 \pi k\). Hence:\[\bf 2x=\frac{ \pi }{ 4 }+ 2 \pi k \implies x = \frac{ \pi }{ 8 }+ \pi k\]

OpenStudy (kenljw):

x = inversecos(1/sprt(2)/2, inversecos(1/sprt(2)/2 + PI/2 x = PI/4, PI/4 + PI/2

OpenStudy (anonymous):

\[2x=\frac{ \pi }{ 4 }\] Solve for x like any other equation.

OpenStudy (anonymous):

Now plug in integer values of 'k' to find all values of 'x' between 0 and 2pi. @pokemonmaster96

OpenStudy (anonymous):

genius solution is much prettier and more general than mine. If you can remember that, then you could apply that for any equation, not just this one.

OpenStudy (anonymous):

I forgot to add, since the angle can actually be pi/8 or 7pi/8 then:\[\bf x=\frac{ \pi }{ 8 }+ \pi k \ \ or \ \ x=\frac{ 7 \pi }{ 8 } + \pi k\]

OpenStudy (anonymous):

And the only values of k that satisfy the inequality are k = 0 and k = 1. PLug them in for k in to both and you'll get 4 solutions. @pokemonmaster96

OpenStudy (anonymous):

Nope. This is cos(2x) not cos(x).

OpenStudy (anonymous):

I got 4 solutions, pi/8, 9pi/8, 7pi/4, and 3pi/4. Right? @genius12

OpenStudy (anonymous):

@pokemonmaster96 The first 2 are right. The other 2 are wrong. Look again at what I wrote 2-3 posts above.

OpenStudy (anonymous):

pi/4 and 5pi/4? If not, then, I am completely lost.

OpenStudy (anonymous):

@genius12 ^^^

OpenStudy (anonymous):

@pokemonmaster96 The other two solutions would be 7pi/8 and 15pi/8.

OpenStudy (kenljw):

2x = inversecos(1/sprt(2), inversecos(1/sprt(2) + PI/2 x = inversecos(1/sprt(2)/2, 2 inversecos(1/sprt(2) + PI/2

OpenStudy (anonymous):

Since, it could be two angles, right? The, pi/8 and 7pi/8. @genius12